OpenStudy (anonymous):

Draw a number line and label it from 0 to 1, numbering every 1/10 increment.

OpenStudy (accessdenied):

How many increments will you have in total? From 0 to 1, labeling every 1/10 point.

OpenStudy (anonymous):

im not sure, i think 1

OpenStudy (accessdenied):

if it helps, 1 is equal to any number divided by itself. 1 = 10/10 So in terms of 1/10, 1 = 10/10 = 1/10 * 10.

OpenStudy (accessdenied):

so we would have 10 such increments on our number line, each being a multiple of 1/10: 1/10, 2/10, 3/10, ... up to 10/10 which is 1.

OpenStudy (anonymous):

oh ok that makes since a little more better, thanks for the help :)

OpenStudy (accessdenied):

you're welcome! :)